OpenStudy (anonymous):

can someone please explain how the integral of 1/(1+2x)^2 = -1/2(2x+1)

OpenStudy (empty):

Yeah, make the substitution: \(1+2x = u\) then differentiate this to get: \(2 = \frac{du}{dx}\) multiply both sides of this by dx to get: \(2dx=du\) Now divide both sides by 2: \(dx = \frac{1}{2}du\) So when you substitute into your integral: \(\int \frac{1}{(1+2x)^2} dx = \int \frac{1}{u^2} \frac{1}{2}du\) Now it should be a little easier to solve from here, then you can substitute back in u=1+2x when you're done.

OpenStudy (anonymous):

omg thank you so much i understand it now

OpenStudy (empty):

Technically you're not allowed to multiply by "dx" like that, but it won't really be a problem until you get to multivariable calculus.

ganeshie8 (ganeshie8):

After doing few problems, it becomes easy to guess these directly notice that the derivative of \(\dfrac{1}{x}\) is \(-\dfrac{1}{x^2}\), so \(-\dfrac{1}{x^2}\) is an antiderivative of \(\dfrac{1}{x}\)

ganeshie8 (ganeshie8):

similarly, if you see that the derivative of \(\dfrac{1}{x+1}\) is \(-\dfrac{1}{(x+1)^2}\), you can say immediately that \(-\dfrac{1}{(x+1)^2}\) is an antiderivative of \(\dfrac{1}{x+1}\)

OpenStudy (anonymous):

that is cool stuff to know this will help out alot

ganeshie8 (ganeshie8):

it is called "advanced guessing", becomes easy with some practice

OpenStudy (empty):

Lol "advanced guessing"

OpenStudy (empty):

Is that really called that or are you just joking around XD

OpenStudy (anonymous):

thats funny

ganeshie8 (ganeshie8):

Haha here is my theory : guessing \(\cos x\) is an antiderivative of \(\sin x\) is just "plain guessing"

ganeshie8 (ganeshie8):

but guessing \(e^{2x}\) is an antiderivative of \(2e^{2x}\) is "advanced guessing" as we're jumping ahead avoiding u substitution

OpenStudy (anonymous):

that is awesome by that theory i have integral of 1/(2x+1)^2 = -1/(2x+1) that is sooo awesome

ganeshie8 (ganeshie8):

i do remember reading that phrase "advanced guessing" from some textbook :)

ganeshie8 (ganeshie8):

Yes! I see that you're getting hang of it, but you will need to be somewhat careful here your guess is a good one, the antiderivative of `1/(2x+1)^2` should look something like `-1/(2x+1) `

ganeshie8 (ganeshie8):

but thats only the first step, to be sure, differentiate `-1/(2x+1) ` and see if you really get ` 1/(2x+1)^2`

OpenStudy (empty):

Yeah, true integrals can be hard but derivatives are easy. So whenever you solve an integral, you should try to check yourself by just differentiating it. It's a useful trick and you'll get better at derivatives in the process so it's kind of like a win-win situation haha.

OpenStudy (anonymous):

ok thats right i see that is important i need to keep that in mind that checking the derivative at the end is important

OpenStudy (anonymous):

thank you guys for all the help, i gotta get to bed

ganeshie8 (ganeshie8):

gnite!